A room has a perimeter of 44 ft. The width is five-sixths of the length. What are the dimensions of the room?
Answer by sonya(5)

Let's say that width=x, and length=y. Then (5/6)y=x. We know that 2(x+y)=44, or 2((5/6)y+(6/6)y)=44. So that means that 2((11/6)y)=44, or (22/6)y=44, or y=44*(22/6). So we have y=12. Since y=12, x=12*(5/6), or x=10.
Check:
2(12+10)=2(22)=44